Determine the net exports and the GDP at market Task 1 - Macroeconomic Accounting The European Commission published the following data for the EU-27 in 2014: Abbildung 1: Use of GDP 7.709,5 billion EUR 2.838,8 billion EUR 202,5 billion EUR 2.605,6 billion EUR 5.801,2 billion EUR 5.452,4 billion EUR Public Consumption Stock changes and net additional valuables Gross capital investment Imports a) Determine the net exports and the GDP at market prices b) Compare the share of private consumption in GDP and the share of public consumption in GDP of the EU-27 with the respective values for Germany. (Data for Germany 2014: Private consumption 1.571,5 billion EUR, public consumption 541,2 billion EUR, GDP: 3006bilon EUR) c) Which impact do the net exports have on the GDP of the EU-27 and the German GDP? (net exports Germany: 163, 2 billion EUR)prices
Answer #1

a) Net exports = Exports - Imports = 5.801,2 - 5.452,4 = 0.348,8 billion EUR

GDP of EU -27 in 2004 = private consumption + government expenditure ( public consumption) + Investment + Net additional values + Net exports

= 7.709,5 + 2.838,8 + 202,5 + 2.605,6 + 0.348,8

= 13.705,2 billion EUR

b) Share of private consumption in the GDP of EU-27 = ( Private consumption / GDP ) * 100 = (7.709,5/13.705,2)*100 = 56.25%

Share of private consumption in the GDP of Germany = (1.571,5/3.006)*100 = 52.53%

So share of private consumption in GDP is more for EU-27 than Germany.

Share of public consumption in the GDP of EU-27 = ( Public consumption / GDP ) * 100 = (2.838,8/13.705,2)*100 = 20.71%

Share of public consumption in the GDP of Germany = ( Public consumption / GDP ) * 100 = (0.541,2/3.006) * 100 = 18%

So share of public consumption in GDP is more for EU-27 than Germany.

c) Share of net exports in the GDP of Germany =( 0.163,2 / 3.006 ) * 100 = 5.429%

So for both , the net exports have a positive impact and the share of net exports in GDP.